OpenStudy (anonymous):

First one to answer gets a medal. What fraction of a circle is 33 1/3percent of a circle? First one to answer gets a medal. What fraction of a circle is 33 1/3percent of a circle? @Mathematics

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Show your work

OpenStudy (anonymous):

33 1/3 divided by 100 = 1/3

OpenStudy (anonymous):

33.3% = .333 ?

OpenStudy (anonymous):

if you have a percent, you shift the decimal two places to the left to get the fraction. Its the definition of a percent.

jimthompson5910 (jim_thompson5910):

33 1/3 = 33+1/3 = 99/3 + 1/3 = 100/3 So 33 1/3 % = 100/3 % Now divide the fraction by 100 to convert to a fraction (100/3)/100 = (100/3)*(1/100) = 100/300 = 1/3 So 100/3 % = 1/3 This means that 33 1/3 % = 1/3

OpenStudy (anonymous):

that's right, but dang youre fast with the script